# MOFFlow: Flow Matching for Structure Prediction of Metal-Organic Frameworks

### v1

@@ Abstract
Metal-organic frameworks (MOFs) are a class of crystalline materials with promising applications in many areas such as carbon capture and drug delivery. In this work, we introduce MOFFlow, the first deep generative model tailored for MOF structure prediction. Existing approaches, including ab initio calculations and even deep generative models, struggle with the complexity of MOF structures due to the large number of atoms in the unit cells. To address this limitation, we propose a novel Riemannian flow matching framework that reduces the dimensionality of the problem by treating the metal nodes and organic linkers as rigid bodies, capitalizing on the inherent modularity of MOFs. By operating in the $SE(3)$ space, MOFFlow effectively captures the roto-translational dynamics of these rigid components in a scalable way. Our experiment demonstrates that MOFFlow accurately predicts MOF structures containing several hundred atoms, significantly outperforming conventional methods and state-of-the-art machine learning baselines while being much faster.
@@ 2 Related work
Flow matching. Flow matching is a simulation-free approach for training continuous normalizing flows, originally introduced by Lipman et al. (2023). Since its introduction, various extensions have been proposed, such as generalization to Riemannian manifolds (Chen & Lipman, 2024) and efficiency improvements through optimal transport (Tong et al., 2024; Pooladian et al., 2023). Due to its flexibility and computational efficiency, flow matching has made notable progress in several related domains, including protein generation (Yim et al., 2023a; b; 2024; Bose et al., 2024), molecular conformation generation (Song et al., 2024), and CSP (Miller et al., 2024).
@@ 3.1 Representation of MOF structures
MOF representation. The 3D crystal structure of a MOF can be represented with the periodic arrangement of the smallest repeating unit called the unit cell. A unit cell containing $N$ atoms can be represented with the tuple $\mathcal{S}=(\bm{X},\bm{a},\bm{\ell})$ where $\bm{X}=[x_{n}]_{n=1}^{N}\in\mathbb{R}^{N\times 3}$ is the atom coordinates, $\bm{a}=[a_{n}]_{n=1}^{N}\in\mathcal{A}^{N}$ is the atom types with $\mathcal{A}$ denoting the set of possible elements, and $\bm{\ell}=(a,b,c,\alpha,\beta,\gamma)\in\mathbb{R}_{+}^{3}\times[0,180]^{3}$ is the lattice parameter that describes the periodicity of the structure (Miller et al., 2024; Luo et al., 2024). In particular, the lattice parameter $\bm{\ell}$ can be transformed into a standard lattice matrix $L\in\mathbb{R}^{3\times 3}$, which defines the infinite crystal structure:
$\{(a_{n}^{\prime},x_{n}^{\prime})|a_{n}^{\prime}=a_{n},x_{n}^{\prime}=x_{n}+Lk,k\in\mathbb{Z}^{3}\}$ (1)
@@
Block-wise representation of MOFs. Here we introduce the blockwise representation of MOFs that decompose a given unit cell into constituent building blocks, i.e., the metal nodes and organic linkers. The blockwise representation is a tuple $\mathcal{S}=(\mathcal{B},\bm{q},\bm{\tau},\bm{\ell})$ where =$\mathcal{B}=\{\mathcal{C}_{m}\}_{m=1}^{M}$ corresponds to $M$ building blocks and $\bm{q}=[q_{m}]_{m=1}^{M},\bm{\tau}=[\tau_{m}]_{m=1}^{M}$ corresponds to set of $M$ building block roto-translations $(q_{m},\tau_{m})\in SE(3)$ Moreover, each block $\mathcal{C}_{m}=(\bm{a}_{m},\bm{Y}_{m})$ has $N_{m}$ atoms with atom types $\bm{a}_{m}=[a_{n}]_{n=1}^{N_{m}}\in\mathcal{A}^{N_{m}}$ and local coordinates $\bm{Y}_{m}=[y_{n}]_{n=1}^{N_{m}}\in\mathbb{R}^{N_{m}\times 3}$. Our main assumption is that the building blocks can be composed by the roto-translations to form the MOF structure, i.e., the atomwise representation $(\bm{X},\bm{a},\bm{\ell})$ can be expressed by the blockwise representation $(\bm{q},\bm{\tau},\mathcal{B},\bm{\ell})$:
$\bm{X}=\operatorname{Concat}(\bm{X}_{1},\ldots,\bm{X}_{M}),\quad\bm{X}_{m}=(q_{m},\tau_{m})\cdot\bm{Y}_{m},$
where $\bm{X}_{m}$ is the result of the roto-translated local coordinate represented by group action $\cdot$. We express the global coordinate $\bm{X}$ as the concatenation $\bm{X}_{m}$’s without loss of generality.
@@ 3.2 Flow matching on Riemannian manifolds
Conditional flow matching on Riemannian manifold. The goal of CNF is to learn a vector field $u_{t}(\cdot)$ that transforms a simple prior distribution $p_{0}$ so that $p_{1}$ closely approximates some target distribution $q$. Given a vector field $u_{t}(\cdot)$ and the corresponding probability paths $\{p_{t}\}_{t\in[0,1]}$, one can train a neural network $v_{t}(x;\theta)$ with the flow matching objective:
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t,p_{t}(x)}[\lVert v_{t}(x;\theta)-u_{t}(x)\rVert_{g}^{2}],$ (3)
where $t\sim\mathcal{U}[0,1]$, $x\sim p_{t}(x)$, and $\lVert\cdot\rVert_{g}$ is the norm induced by the metric $g$. However, the flow matching objective lacks analytic form of $u_{t}(x)$ that transforms the prior $p_{0}$ into the target $q$. The key insight of conditional flow matching objective is to instead learn a conditional vector field $u_{t}(x|x_{1})$ for a data point $x_{1}$ defined as follows:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,p_{1}(x),p_{t}(x|x_{1})}[\lVert v_{t}(x;\theta)-u_{t}(x|x_{1})\rVert_{g}^{2}],$ (4)
where we let $p_{0}(x|x_{1})=p_{0}(x)$ and $p_{1}(x|x_{1})\approx\delta(x-x_{1})$ with $\delta(\cdot)$ being the Dirac distribution. The key idea of conditional flow matching is that one can derive the conditional vector field $u_{t}(x|x_{1})$ that marginalizes over data points $x_{1}\sim q$ accordingly to induce the vector field $u_{t}(x)$ transforming the prior $p_{0}$ into the desired distribution $q$. To construct $u_{t}(x|x_{1})$, Chen & Lipman (2024) proposes defining conditional flow $x_{t}=\phi_{t}(x_{0}|x_{1})$ from the geodesic path (minimum length curve) connecting two points $x_{0},x_{1}\in\mathcal{M}$ by $x_{t}=\exp_{x_{0}}(t\log_{x_{0}}(x_{1}))$ for $t\in[0,1]$, where $\exp_{x}$ and $\log_{x}$ are the exponential and logarithmic map at point $x\in\mathcal{M}$, respectively. Then the desired conditional vector field can be derived as the time derivative, i.e., $u(x_{t}|x_{1})=\frac{d}{dt}\phi_{t}(x_{0}|x_{1})$.
@@ 4.1 Construction of local coordinates
Here, $C(\bm{X})$ denotes subtraction of the centroid and $\mathcal{R}(\bm{a},\bm{X})$ denotes rotation to align the building block with the PCA axes whose sign is fixed by a reference vector, following Gao & Günnemann (2022). See Appendix D for more details.
@@ 4.2 Flow matching for MOF structure prediction
Metric for $SE(3)$. Following Yim et al. (2023b), we treat $SO(3)$ and $\mathbb{R}^{3}$ independently by defining an additive metric on $SE(3)$ as
$\langle(q,\tau),(q^{\prime},\tau^{\prime})\rangle_{SE(3)}=\langle q,q^{\prime}\rangle_{SO(3)}+\langle\tau,\tau^{\prime}\rangle_{\mathbb{R}^{3}}.$ (7)
Here, $\langle q,q^{\prime}\rangle_{SO(3)}$ and $\langle\tau,\tau^{\prime}\rangle_{\mathbb{R}^{3}}$ are inner products defined as $\langle q,q^{\prime}\rangle_{SO(3)}=\text{tr}(qq^{\prime\top})/2$ and $\langle\tau,\tau^{\prime}\rangle_{\mathbb{R}^{3}}=\tau^{\top}\tau^{\prime}$ for $q,q^{\prime}\in\mathfrak{so}(3)$ with $\mathfrak{so}(3)$ denoting the Lie algebra of $SO(3)$ and $\tau,\tau^{\prime}\in\mathbb{R}^{3}$.
@@
$\mathcal{L}(\theta)=$ $\mathbb{E}_{\mathcal{M}^{(1)}\sim\mathcal{D}}\mathbb{E}_{t\in\mathcal{U}(0,1)}\big{[}\lambda_{1}\lVert\bm{\hat{q}}_{1}-\bm{q}_{1}\rVert_{SO(3)}+\lambda_{2}\lVert\bm{\hat{\tau}}_{1}-\bm{\tau}_{1}\rVert_{\mathbb{R}^{3}}+\lambda_{3}\lVert\hat{\bm{\ell}}_{1}-\bm{\ell}_{1}\rVert_{\mathbb{R}^{3}}\big{]},$ (10)
@@ 4.3 Model Architecture
(a) Atom-level update layer
(b) Block-level update modules
Figure 3: Overview of our neural network architecture. (3(a)) The atom-level update layers encode building block representations into embeddings at the atomic resolution. (3(b)) The block-level modules iteratively update the roto-translations $(\bm{q},\bm{\tau})$, block features $\bm{H}$, pairwise features $\bm{Z}$, and lattice parameter $\bm{\ell}$ to predict the clean data $(\hat{\bm{q}}_{1},\hat{\bm{\tau}}_{1},\hat{\bm{\ell}}_{1})$.
@@
Block-level update modules. Each layer of the update module (Figure 3(b)) iteratively updates its prediction of $(\bm{q},\bm{\tau})\in{SE(3)^{M}}$ and $\ell$ along with the block features $\bm{H}=[h_{m}]_{m=1}^{M}\in\mathbb{R}^{M\times d_{h}}$ and the pairwise features $\bm{Z}=[z_{mm^{\prime}}]_{m,m^{\prime}=1}^{M}\in\mathbb{R}^{M\times M\times d_{z}}$. The predictions are initialized by the intermediate flow matching output $\bm{q}_{t},\bm{\tau}_{t},\ell_{t}$, the node features are initialized from the atom-level update layers, and the edge features are initialized as follows:
$z_{mm^{\prime}}=\phi_{z}(h_{m},h_{m^{\prime}},\operatorname{dgram}({\lVert\tau_{m}-\tau_{m^{\prime}}\rVert}_{2}),\operatorname{dgram}({\lVert\hat{\tau}_{m}-\hat{\tau}_{m^{\prime}}\rVert}_{2})),$ (12)
@@
$\bm{H}^{\prime}$ $=\operatorname{NodeUpdate}(\bm{H},\bm{Z},\bm{q},\bm{\tau},\ell),\quad\bm{Z}^{\prime}=\operatorname{EdgeUpdate}(\bm{Z},\bm{H}^{\prime}),$ (13)
$(\bm{q}^{\prime},\bm{\tau}^{\prime})$ $=\operatorname{BackboneUpdate}(\bm{q},\bm{\tau},\bm{H}^{\prime}),\quad\ell^{\prime}=\operatorname{LatticeUpdate}(\ell,\bm{H}^{\prime}),$ (14)
where $\bm{q}^{\prime},\bm{\tau}^{\prime},\bm{H}^{\prime},\bm{Z}^{\prime},\ell^{\prime}$ are the updated predictions and the features. Importantly, $\operatorname{NodeUpdate}$ operator consists of the newly designed $\operatorname{MOFAttention}$ module followed by pre-layer normalization $\operatorname{Transformers}$ (Xiong et al., 2020) and MLP with residual connections in between. The $\operatorname{EdgeUpdate}$, $\operatorname{BackboneUpdate}$ modules are implemented as in Yim et al. (2023b) and $\operatorname{LatticeUpdate}$ is identity function for the lattice parameter except for the last layer. At the final layer, the lattice parameter is predicted from mean pooling of the block features followed by an MLP. Complete details of each update module are in Appendix E.
@@ 5.1 Structure prediction
Implementation. Both RS and EA use CHGnet (Deng et al., 2023) for structure optimization. We generate 20 samples for RS, while EA starts with 5 initial, 20 populations, and 20 generations. Due to the high computational cost for large crystals, generating more samples was not feasible. For DiffCSP, we follow the hyperparameters from Jiao et al. (2024a) and train for 200 epochs. Our method uses an AdamW optimizer (Loshchilov, 2017) with a learning rate of $10^{-4}$ and $\beta=(0.9,0.98)$. We use a maximum batch size of 160 and run inference with 50 integration steps. We generate $1$ and $5$ samples for DiffCSP and our method. Implementation details are in Appendix C.
@@
Table 1: Structure prediction accuracy. We evaluate structure prediction performance across optimization-based methods (RS, EA), a deep generative model (DiffCSP), and our method, MOFFlow. Due to computational constraints, RS and EA were tested on 100 and 15 samples, respectively, while DiffCSP and MOFFlow were evaluated on the full test set of 30,880 structures. MR refers to the match rate, and RMSE to the root mean squared error; - indicates no match. $\operatorname{stol}$ represents the site tolerance for matching criteria. The reported time is the average per structure. MOFFlow outperforms all baselines in MR, RMSE, and generation time.
@@
Figure 4: Visualization of the predicted MOF structures. We select structures from the 20 candidates with the lowest RMSE. The lattice is scaled to represent the relative size between generated structures. Our model accurately generates high-quality MOF structure predictions with accurate atomic positions and lattice configuration.

### v2

@@ Abstract
Metal-organic frameworks (MOFs) are a class of crystalline materials with promising applications in many areas such as carbon capture and drug delivery. In this work, we introduce MOFFlow, the first deep generative model tailored for MOF structure prediction. Existing approaches, including ab initio calculations and even deep generative models, struggle with the complexity of MOF structures due to the large number of atoms in the unit cells. To address this limitation, we propose a novel Riemannian flow matching framework that reduces the dimensionality of the problem by treating the metal nodes and organic linkers as rigid bodies, capitalizing on the inherent modularity of MOFs. By operating in the $SE(3)$ space, MOFFlow effectively captures the roto-translational dynamics of these rigid components in a scalable way. Our experiment demonstrates that MOFFlow accurately predicts MOF structures containing several hundred atoms, significantly outperforming conventional methods and state-of-the-art machine learning baselines while being much faster. Code available at https://github.com/nayoung10/MOFFlow.
@@ 2 Related work
Flow matching. Flow matching is a simulation-free approach for training continuous normalizing flows (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023). Since its introduction, various extensions have been proposed, such as generalization to Riemannian manifolds (Chen & Lipman, 2024) and efficiency improvements through optimal transport (Tong et al., 2024; Pooladian et al., 2023). Due to its flexibility and computational efficiency, flow matching has made notable progress in several related domains, including protein generation (Yim et al., 2023a; b; 2024; Bose et al., 2024), molecular conformation generation (Song et al., 2024), and CSP (Miller et al., 2024).
@@ 3.1 Representation of MOF structures
MOF representation. The 3D crystal structure of a MOF can be represented with the periodic arrangement of the smallest repeating unit called the unit cell. A unit cell containing $N$ atoms can be represented with the tuple $\mathcal{S}=(\bm{X},\bm{a},\bm{\ell})$ where $\bm{X}=[x_{n}]_{n=1}^{N}\in\mathbb{R}^{N\times 3}$ is the atom coordinates, $\bm{a}=[a_{n}]_{n=1}^{N}\in\mathcal{A}^{N}$ is the atom types with $\mathcal{A}$ denoting the set of possible elements, and $\bm{\ell}=(a,b,c,\alpha,\beta,\gamma)\in\mathbb{R}_{+}^{3}\times[0,180]^{3}$ is the lattice parameter that describes the periodicity of the structure (Miller et al., 2024; Luo et al., 2024). In particular, the lattice parameter $\bm{\ell}$ can be transformed into a standard lattice matrix $L=(\bm{l}_{1},\bm{l}_{2},\bm{l}_{3})\in\mathbb{R}^{3\times 3}$, which defines the infinite crystal structure:
$\{(a_{n}^{\prime},x_{n}^{\prime})|a_{n}^{\prime}=a_{n},x_{n}^{\prime}=x_{n}+kL^{\top},k\in\mathbb{Z}^{1\times 3}\}$ (1)
@@
Block-wise representation of MOFs. Here we introduce the blockwise representation of MOFs that decompose a given unit cell into constituent building blocks, i.e., the metal nodes and organic linkers. The blockwise representation is a tuple $\mathcal{S}=(\mathcal{B},\bm{q},\bm{\tau},\bm{\ell})$ where =$\mathcal{B}=\{\mathcal{C}_{m}\}_{m=1}^{M}$ corresponds to $M$ building blocks and $\bm{q}=[q_{m}]_{m=1}^{M},\bm{\tau}=[\tau_{m}]_{m=1}^{M}$ corresponds to set of $M$ building block roto-translations $(q_{m},\tau_{m})\in SE(3)$ Moreover, each block $\mathcal{C}_{m}=(\bm{a}_{m},\bm{Y}_{m})$ has $N_{m}$ atoms with atom types $\bm{a}_{m}=[a_{n}]_{n=1}^{N_{m}}\in\mathcal{A}^{N_{m}}$ and local coordinates $\bm{Y}_{m}=[y_{n}]_{n=1}^{N_{m}}\in\mathbb{R}^{N_{m}\times 3}$ (defined in Section 4.1). Our main assumption is that the building blocks can be composed by the roto-translations to form the MOF structure, i.e., the atomwise representation $(\bm{X},\bm{a},\bm{\ell})$ can be expressed by the blockwise representation $(\bm{q},\bm{\tau},\mathcal{B},\bm{\ell})$ as $\bm{X}=\operatorname{Concat}(\bm{X}_{1},\ldots,\bm{X}_{M}),\bm{X}_{m}=(q_{m},\tau_{m})\cdot\bm{Y}_{m},$ where $\bm{X}_{m}$ is the result of the roto-translated local coordinate represented by group action $\cdot$. We express the global coordinate $\bm{X}$ as the concatenation $\bm{X}_{m}$’s without loss of generality.
@@ 3.2 Flow matching on Riemannian manifolds
Conditional flow matching on Riemannian manifold. The goal of CNF is to learn a vector field $u_{t}(\cdot)$ that transforms a simple prior distribution $p_{0}$ so that $p_{1}$ closely approximates some target distribution $q$. Given a vector field $u_{t}(\cdot)$ and the corresponding probability paths $\{p_{t}\}_{t\in[0,1]}$, one can train a neural network $v_{t}(x;\theta)$ with the flow matching objective $\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t,p_{t}(x)}[\lVert v_{t}(x;\theta)-u_{t}(x)\rVert_{g}^{2}],$ where $t\sim\mathcal{U}[0,1]$, $x\sim p_{t}(x)$, and $\lVert\cdot\rVert_{g}$ is the norm induced by the metric $g$. However, the flow matching objective lacks analytic form of $u_{t}(x)$ that transforms the prior $p_{0}$ into the target $q$. The key insight of conditional flow matching objective is to instead learn a conditional vector field $u_{t}(x|x_{1})$ for a data point $x_{1}$ defined as $\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,p_{1}(x),p_{t}(x|x_{1})}[\lVert v_{t}(x;\theta)-u_{t}(x|x_{1})\rVert_{g}^{2}],$ where we let $p_{0}(x|x_{1})=p_{0}(x)$ and $p_{1}(x|x_{1})\approx\delta(x-x_{1})$ with $\delta(\cdot)$ being the Dirac distribution. The key idea of conditional flow matching is that one can derive the conditional vector field $u_{t}(x|x_{1})$ that marginalizes over data points $x_{1}\sim q$ accordingly to induce the vector field $u_{t}(x)$ transforming the prior $p_{0}$ into the desired distribution $q$. To construct $u_{t}(x|x_{1})$, Chen & Lipman (2024) proposes defining conditional flow $x_{t}=\phi_{t}(x_{0}|x_{1})$ from the geodesic path (minimum length curve) connecting two points $x_{0},x_{1}\in\mathcal{M}$ by $x_{t}=\exp_{x_{0}}(t\log_{x_{0}}(x_{1}))$ for $t\in[0,1]$, where $\exp_{x}$ and $\log_{x}$ are the exponential and logarithmic map at point $x\in\mathcal{M}$, respectively. Then the desired conditional vector field can be derived as the time derivative, i.e., $u(x_{t}|x_{1})=\frac{d}{dt}\phi_{t}(x_{0}|x_{1})$.
@@ 4.1 Construction of local coordinates
Here, $C(\bm{X})$ denotes subtraction of the centroid and $\mathcal{R}(\bm{a},\bm{X})$ denotes rotation to align the building block with the PCA axes whose sign is fixed by a reference vector, following Gao & Günnemann (2022). See Appendix E for more details.
@@ 4.2 Flow matching for MOF structure prediction
Metric for $SE(3)$. Following Yim et al. (2023b), we treat $SO(3)$ and $\mathbb{R}^{3}$ independently by defining an additive metric on $SE(3)$ as $\langle(q,\tau),(q^{\prime},\tau^{\prime})\rangle_{SE(3)}=\langle q,q^{\prime}\rangle_{SO(3)}+\langle\tau,\tau^{\prime}\rangle_{\mathbb{R}^{3}}.$ Here, $\langle q,q^{\prime}\rangle_{SO(3)}$ and $\langle\tau,\tau^{\prime}\rangle_{\mathbb{R}^{3}}$ are inner products defined as $\langle q,q^{\prime}\rangle_{SO(3)}=\text{tr}(qq^{\prime\top})/2$ and $\langle\tau,\tau^{\prime}\rangle_{\mathbb{R}^{3}}=\tau^{\top}\tau^{\prime}$ for $q,q^{\prime}\in\mathfrak{so}(3)$ with $\mathfrak{so}(3)$ denoting the Lie algebra of $SO(3)$ and $\tau,\tau^{\prime}\in\mathbb{R}^{3}$.
@@
$\mathcal{L}(\theta)=\mathbb{E}_{\mathcal{S}^{(1)}\sim\mathcal{D}}\mathbb{E}_{t\sim\mathcal{U}(0,1)}\Big{[}\frac{1}{(1-t)^{2}}\Big{(}\lambda_{1}$ $\lVert\log_{\bf{q}^{(t)}}(\bm{\hat{q}}_{1})-\log_{\bf{q}^{(t)}}(\bm{q}_{1})\rVert_{SO(3)}^{2}$
$+\lambda_{2}$ $\lVert\bm{\hat{\tau}}_{1}-\bm{\tau}_{1}\rVert_{\mathbb{R}^{3}}^{2}+\lambda_{3}\lVert\hat{\bm{\ell}}_{1}-\bm{\ell}_{1}\rVert_{\mathbb{R}^{3}}^{2}\Big{)}\Big{]},$ (7)
@@ 4.3 Model Architecture
Figure 3: Overview of our neural network architecture. Our architecture follows a hierarchical structure, starting with atom-level update layers that encode building block representations into atomic-resolution embeddings. These are followed by block-level update layers, which iteratively refine the roto-translations $(\bm{q},\bm{\tau})$, block features $\bm{H}$, pairwise features $\bm{Z}$, and lattice parameters $\bm{\ell}$. The final output is a prediction of the clean data $(\hat{\bm{q}}_{1},\hat{\bm{\tau}}_{1},\hat{\bm{\ell}}_{1})$.
@@ 5.1 Structure prediction
Implementation. Both RS and EA use CHGnet (Deng et al., 2023) for structure optimization. We generate 20 samples for RS, while EA starts with 5 initial, 20 populations, and 20 generations. Due to the high computational cost for large crystals, generating more samples was not feasible. For DiffCSP, we follow the hyperparameters from Jiao et al. (2024a) and train for 200 epochs. Our method uses an AdamW optimizer (Loshchilov, 2017) with a learning rate of $10^{-4}$ and $\beta=(0.9,0.98)$. We use a maximum batch size of 160 and run inference with 50 integration steps (see Appendix G for analysis on integration steps). We generate $1$ and $5$ samples for DiffCSP and our method. Implementation details are in Appendix C.
@@
Table 1: Structure prediction accuracy. We compare optimization-based methods (RS, EA), a deep generative model (DiffCSP), and MOFFlow. Due to computational constraints, RS and EA were tested on 100 and 15 samples, respectively, while DiffCSP and MOFFlow were evaluated on the full test set (30,880 structures). MR is the match rate and RMSE is the root mean squared error; - indicates no match. $\operatorname{stol}$ is the site tolerance for matching criteria. The reported time is the average per structure. MOFFlow outperforms all baselines in MR, RMSE, and generation time.
@@
Figure 4: Visualization of the predicted MOF structures. We select structures from the 20 candidates with the lowest RMSE. The lattice is scaled to reflect the relative sizes. MOFFlow accurately generates high-quality predictions with accurate atomic positions and lattice configuration.